OpenStudy (anonymous):

simplify 4x^6-8x^3+18=0

OpenStudy (anonymous):

any common factors?

OpenStudy (tkhunny):

?? "Simplify", or "Solve"? To Solve, it may be important to notice that it is QUADRATIC. No, really! \(x^{3} = \dfrac{-b\pm\sqrt{b^{2} - 4ac}}{2a}\) If it really is "simplify", listen to aacehm! Find those common factors.

OpenStudy (anonymous):

@tkhunny is right, if you set u = x^3, then it turns into a quadratic \[4u^{2}-8u+18\]